Two alloys contain silver and copper in the ratios 3:1 and 5:3 respectively.The alloys are mixed to get a third alloy.The possible ratio of silver to copper in the third alloy is?
Answer by vasumathi(46)

Solution :The original ratio in the first alloy
silver : copper = 3:1 (if 3 parts silver then 1 part copper so total = 3+1 = 4)
the original ratio in the second alloy
silver : copper = 5:3(if 5 parts silver then 3 parts copper so total = 5+3 = 8)
If we mix both these alloys..
there are 3 +5 = 8 parts silver and 1+3= 4 parts silver
so the ratio in the third alloy is...
silver : copper= 8:4 = 2:1 (on diving each term of the ratio by 2)
So the final answer is 2:1